OpenStudy (anonymous):

i dont get this problem?? i seriously cant figure out how to do this stuff. please explain your answer please please please :):):) 5x-7=7x-17 3x+9=44-2x 3x+6=x=2 please and thank you!!!! :)

OpenStudy (anonymous):

oops the last one is 3x+6=x+2

OpenStudy (anonymous):

okkkkkkk

OpenStudy (anonymous):

??

OpenStudy (anonymous):

hold on let me explain this correctly all at once xD

OpenStudy (anonymous):

haha ok :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

????

OpenStudy (anonymous):

you still there??

OpenStudy (anonymous):

5x-7=7x-17 The goal is to get x on one side of the equation by itself. First you subtract 7x from both sides and end up with -2x - 7 = -17 Then you add 7 on both sides -2x = -10 Then you divide both sides by two -2 / -2 cancels out so that just leaves x -10 / -2 is positive 5 because a negative and a negative equals a positive. That leaves you with the equation x = 5

OpenStudy (anonymous):

i still sont get it!!! like ill get it hen i try a different equation and i forget or something or i forget which side i need to subract!! im so sorry

OpenStudy (anonymous):

I understand, it's okay. I still get confused on which side to subtract from! hold on :)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

http://www.wikihow.com/Solve-a-Simple-Linear-Equation Tell me if this helps your understanding :)

OpenStudy (anonymous):

@miraclemilktea2 are you still there??

OpenStudy (anonymous):

If not, try this http://www.purplemath.com/modules/solvelin.htm :3

OpenStudy (anonymous):

ugh i still dont gt it :(:(:(

OpenStudy (anonymous):

hmm do you know what like terms are?

OpenStudy (anonymous):

like 2x and 3x are like terms because they both have x 9y and -2y are like terms because they both have y and 8 and -60 are like terms because they're both regular numbers

OpenStudy (anonymous):

Basically in an equation you want to know what the like terms are first. for example; 5x + 2 = 3x + 1 5x and 3x are like terms 2 and 1 are also like terms

OpenStudy (anonymous):

The rue of any equation is that whatever you do to one side, you must do to the other.

OpenStudy (anonymous):

rule*

OpenStudy (anonymous):

and you want x or your variable alone by itself on one side

OpenStudy (anonymous):

so 5x + 2 = 3x + 1 since we have positive 3 in order to cancel it out we need to subtract 3 (to make it 0) but since whatever we do to one side we have to do to the other we must also subtract 3 from 5x which leaves us with 2x so 2x + 2 = 1 getting it so far?